OpenStudy (anonymous):

x^2 + xy = 1, xy+y^2 = 3, solve for x and y

OpenStudy (anonymous):

Since you have xy in each equation they must be the same result. Solve the first and second equations for xy

OpenStudy (anonymous):

xy = 1 - x^2 and xy = 3 - y^2 1 - x^2 = 3 - y^2

OpenStudy (anonymous):

Solve for y: subtract 3 from both sides -2 - x^2 = -y^2 take the opposite of both sides 2 + x^2 = y^2 square root both sides sqrt(x^2 + 2) = y y = sqrt(x^2 + 2)

OpenStudy (anonymous):

Solve for x: 1 - x^2 = 3 - y^2 subtract 1 from both sides -x^2 = 2 - y^2 take the opposite of both sides x^2 = y^2 - 2 square root both sides x = sqrt(y^2 - 2)

OpenStudy (anonymous):

is there another way to solve this?

OpenStudy (anonymous):

i think the results should be real numbers

OpenStudy (anonymous):

Come on, we can be more elegant than this, guys: Add them both: x^2 + 2xy + y^2 = 4 (x+y)^2 = 4 => x+y = ±2

OpenStudy (anonymous):

Case 1: x = 2-y gives (2-y)y = 3-y^2 => y = 3/2, x = 1/2 Case 2 x = -(2+y) gives -y(2+y) = 3-y^2 gives y = -3/2, x = -1/2

OpenStudy (anonymous):

That's how we do it at Cambridge, featheres!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome, kiddo.

OpenStudy (anonymous):

x=1/2,x=-1/2 and y=3/2 and y=-3/2 are the answers x(x+y)=1 and y(x+y)=3 implies x/y=1/3 and 3x=y .PLug in this in x^2+xy=1,x^2+3x^2=1,4x^2=1, x^2=1,x=1/2 or -1/2 now we have y3x=3(1/2)=3/2 and -3/2